OpenStudy (anonymous):

Solve by taking the square roots of both sides. (x + 4)2 - 5 = 2 FOR MY ANSWER I GOT X=4+√ 7, X=-4+√ 7 BUT MY TEACHER SAID IM WRONG

OpenStudy (anonymous):

go to www.purplemath.com/modules/solvquad2.htm

OpenStudy (anonymous):

it has videos to help

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \it (x+4)^2-5=2 }\) 1) add 5 to both sides 2) take the square root of both sides 3) subtract 4 from both sides. you are correct, but one more addition... \(\Huge\color{blue}{ \it x=-4±\sqrt{7 } }\) 4 is always negative, 7 is ±.

OpenStudy (solomonzelman):

Do you see hwy the teachers say that you are wrong ? They are correct.... -:(

OpenStudy (solomonzelman):

actually it's only 1 teacher.

OpenStudy (solomonzelman):

Good luck !

OpenStudy (anonymous):

try the website ambbiiee

OpenStudy (anonymous):

X=3,X=-11 @SolomonZelman

OpenStudy (solomonzelman):

Noo ! you are completing the square. @Ambbiiee can we start over, b/c you seem to lose me overhere.

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \rm (x + 4)^2 - 5 = 2 }\) @Ambbiiee , show me what you get after you add 5 to both sides please.

OpenStudy (anonymous):

SEARCH ON GOOGLE X+4 IT WILL SHOW A GRAPH

OpenStudy (anonymous):

 @SolomonZelman

OpenStudy (anonymous):

try the website ambbiiee

OpenStudy (solomonzelman):

I got disconnected, sorry -:( \(\Huge\color{blue}{ \sf (x+4)^2=7 }\) yes, that's correct. Now, square root both sides,